# Nearly geodesic surfaces are filling

### v2

@@ 3.4 The dichotomy of embedded surfaces in a hyperbolic $3$-manifold
3.5 A filling geodesic surface is strongly filling
Recall that a homotopy class $[S]$ of surfaces in a hyperbolic manifold $M$ separates $p,q\in\partial\widetilde{M}$ if any $S^{\prime}\in[S]$ has a cover $\widetilde{S}$ such that $\partial\widetilde{S}$ separate $p,q\in\partial\widetilde{M}$. Contractible surfaces in $M$ can be homotoped to be filling. However, [RS99, Lemma 2.1] proved that if a Fuchsian surface $S$ is filling, its homotopy class $[S]$ is strongly filling. We present a slightly simpler proof.
Lemma 3.18.
Proof.
$M-S=\cup_{i=1}^{n}B_{i},\textnormal{ where }B_{i}\textnormal{ is open and contractible},$
[…]
Thus $\gamma\cap\partial B_{i}\neq\emptyset$ for some $i$. If $\gamma$ transversally intersects the interior of some polyhedron $F_{j}$ transversally, by lifting, we find a plane which separates $\partial\tilde{\gamma}$.
The idea of the proof also shows that a totally geodesic hypersurface $S\hookrightarrow M$ is filling if and only if every component of $\widetilde{M}-\pi^{-1}(S)$ is bounded.
Remark 3.19.
The argument above relies on the analysis of the intersection of a geodesic $\gamma$ and a geodesic face $\subset$ a totally geodesic plane $P$ and the dichotomy that either $\tilde{\gamma}\subset\widetilde{P}$ or $\tilde{\gamma}\pitchfork\widetilde{P}$. This gives a geometric criterion for a homotopy class of totally geodesic surfaces to be filling in certain aspherical manifolds. See Section 4.1 for filling homotopy classes of surfaces in Euclidean and $\mathbb{H}^{2}\times\mathbb{R}$ 33-manifolds.
3.6 Cubulations and one-orbit of hyperplanes
Throughout this subsection, let $X$ be a CAT(0) cube complex. We define an equivalence relation on the edges of $X$ generated by identifying opposite edges of some square in $X$. Given an equivalence class $[e]$ of edges, the hyperplane $w$ dual to $[e]$ is the collection of midcubes which intersect edges in $[e]$. The complement XwX-w consists of two halfspaces. Let $G$ be a finitely generated group with a Cayley graph $\Delta$. A subgroup $H\leq G$ is codimension-$1$ if it has a finite neighborhood $N_{r}(H)$ such that $\Delta-N_{r}(H)$ contains at least two $H$-orbits of deep components, which do not lie in any $N_{s}(H)$. A CAT(0) cube complex $X$ is essential, if for each hyperplane $w$, each of the associated halfspaces contains points in $X$ arbitrarily far from $w$. If a group $G$ acts by automorphism on $X$, the action is essential if for any point $x$ in the zero-skeleton of $X$, and each hyperplane $w$, each of the associated halfspaces contains points in $G\cdot x$ arbitrarily far from $w$. The cube complex $X$ is hyperplane-essential if each hyperplane $w$, regarded itself as a CAT(0) cube complex, is essential. The action of $G$ on $X$ is hyperplane-essential if each hyperplane $w$ has the property that the stabilizer of $w$ acts essentially on $w$.
For more backgrounds and basic definitions of cube complex and its connection to $3$-manifolds, we refer the readers to [AFW15, FH21, BW12] and the references therein. Sageev [Sag95] introduced a method to construct a proper action on a CAT(0) cube complex, which elucidates many structures of the groups.
Since a strongly filling surface $[S]$ separates any two points on the Gromov boundary of $G$, one can apply Bergeron-Wise boundary criterion [BW12] to construct a proper action on a CAT(0) cube complex $X$. The singleton $\{[S]\}$ satisfies the assumption of [BW12, Theorem 1.4], and thus its fundamental group $\pi_{1}(S)\rightarrow H\leq\pi_{1}(M)$ is a quasiconvex codimension-$1$ subgroup such that $G$ acts properly and cocompactly on the dual CAT(0) cube complex. The orbits of the hyperplanes correspond to the image $H$ under the natural action by $\pi_{1}(M)$. The stabilizers of the hyperplanes correspond to conjugation of $H\cong\pi_{1}(S)$.
As a consequence of Theorem 1.1, we have the following.
Corollary 3.20.
Let $G$ be the fundamental group of a closed hyperbolic $3$-manifold or a closed arithmetic hyperbolic manifold with a totally geodesic hypersurface. There exist infinitely many CAT(0) cube complexes $X$ on which $G$ acts properly, cocompactly, essentially, and hyperplane-essentially with one orbit of hyperplanes whose stabilizer is a hypersurface subgroup.
Since non commensurable strongly filling surfaces have different limit sets on $\partial G$, they give rise to cubulations that are not equivariantly isometric. This partially answers a question of Fioravanti-Hagen on [FH21, p. 879]. For any Gromov-hyperbolic group $G$ that admits a proper, cocompact action on a CAT(0) cubed complex, Fioravanti-Hagen [FH21] establish the existence of a CAT(0) cube complex $X$ on which $G$ acts properly, cocompactly, and essentially with a single orbit of hyperplanes, or acts properly, cocompactly, essentially, and hyperplane-essentially. Moreover, if we apply their theorems to the case where $G=\pi_{1}(M)$ for $M$ a closed hyperbolic $3$-manifold, it is not clear that their hyperplane-stabilizer is a surface subgroups [^2] M. Hagen informed the author that their construction applied to a filling collection of Fuchsian surfaces should give hyperplane-stabilizers that are surface subgroups. This is because their cubulation comes from taking a filling collection $A$ of finitely many surfaces constructed by [KM12] and then applying the boundary criterion [BW12]. Surfaces in $A$ may coincide on pants instead of just intersecting at closed curves. Thus it is not clear their hyperplane-stabilizers are surface subgroups. In Kahn-Wright’s construction [KW21], one uses two copies of each good pants. See also 3.16.
Fioravanti-Hagen remark on [FH21, p.880] that for certain $G$ and $X$ (including our cases), an arbitrarily small deformation of one of the original walls suffices to obtain a proper action with a single orbit of hyperplanes. It is interesting to note that using completely different techniques, i.e., homogeneous dynamics and rigidity results in [Rat91, MS95, Sha91a], we establish a similar result in our context. In particular, by Theorem 1.6, asymptotically Fuchsian surfaces are dense in $\mathcal{G}_{2}M$ and are eventually filling by Theorem 1.1. Thus in an arbitrarily small neighborhood of a geometric circle at $\partial\widetilde{M}$, there exists the limit set of a strongly filling surface.
Corollary 3.21.
Let $M$ be a closed hyperbolic $3$-manifold or a closed arithmetic hyperbolic manifold with a totally geodesic hypersurface. The collection of strongly filling hypersurfaces is ubiquitous.
@@ 3.7 Least area non-Fuchsian surfaces fail to minimize intersection numbers
Let $M$ be a hyperbolic $3$-manifold. Then for any least area non geodesic surface $S$, there exists a closed geodesic $\gamma$ such that the intersection number $|S\cap\gamma|$ is strictly larger than the geometric intersection number $S]\cdot[\gamma]$.
@@
Take any non geodesic least area surface $S$ of genus $g$ in $M$. Since minimal surfaces in $3$-manifolds are smoothly immersed by 3.10 and non transverse self-intersections are isolated by [FHS83, Lemma 1.4], there is a point $p\in S$ not in the self-intersection locus of $S$. There exists a tangent plane at $p$ and also a totally geodesic plane $\Pi\subset M$ tangent to $S$ at $p$.
@@
We can draw two small open balls $B_{1},B_{2}$ in the two parts that are away from the self-intersection locus. Here we also require that $B_{1},B_{2}$ are supported in a small neighborhood $U$ of $p$, so that $U$ is away from the self-intersection locus. See Figure 3. The set of geodesics which pass through $B_{1}$ and $B_{2}$ in $M$ contains an open subset. By [KH95, Theorem 17.6.2, Corollary 6.4.20], the collection of closed geodesics is a dense subset of the unit tangent bundle of $M$. Hence, there is a closed geodesic $\gamma$ in $M$ which passes through $B_{1}$ and $B_{2}$. Then the geodesic $\gamma$ intersects the least area surface $S$ at two points which are non essential intersections, and can be removed after applying a homotopy to $\gamma$. ∎
@@ 4 From filling surfaces to the topology and geometry of the ambient manifold
Nathan M. Dunfield first suggested the existence of a filling $\pi_{1}$-injective surfaces in non hyperbolic manifolds, such as the Hantzsche-Wendt manifold. In this section, we will classify closed orientable geometric $3$-manifolds which contain a filling $\pi_{1}$-injective surface (Theorem 1.10) and prove the isoperimetric inequality between a filling surface and $M$ (Theorem 1.11).
@@
Let $M$ be a Seifert fiber space with a geometric structure. Let $S$ be a closed surface not equal to $S^{2},P^{2}$, and let $f:S\rightarrow M$ be a minimal $\pi_{1}$-injective immersion. Then $F$ is either vertical or horizontal.
@@
Scott [Sco78] shows that the fundamental group of a Seifert fibered manifold $M$ is LERF. Thus to understand the complementary regions to an immersed $\pi_{1}$-injective surface $S$ in $M$, we can pass to a finite cover $N$ of $M$ so that the preimage of $S$ in $N$ is a finite union of embedded incompressible surfaces that are invariant under the covering group $C$. 3.5 implies that it suffices to show that no such collection of embedded incompressible surfaces fills $N$.
A closed $3$-manifold with $S^{2}\times\mathbb{R}$ geometry is finitely covered by $S^{2}\times S^{1}$. Then $S^{2}\times\theta$ is the only incompressible surface up to homotopy, and any finite union of these surfaces is non-filling.
@@
    the Hantzsche–Wendt manifold.
@@
The $3$-torus has no filling surfaces, since every horizontal surface is a torus whose complement is $\mathbb{T}^{2}\times I$. Using ideas in 3.18 and 3.2, we prove
@@
Readers can compare 4.3 with 3.18.
For $i=2,3,6$, $C_{i}$ denotes the covering group corresponding to the covering $M_{1}\rightarrow M_{i}$. For $i=4,5$, $C_{i}$ refers to the group corresponding to the covering from hexagonal $\mathbb{T}^{3}$ (the hexagonal fundamental domain with opposite faces glued to each other) to $M_{i}$.
@@
The last one is the Hantzsche-Wendt manifold $M_{6}$, which has a $4$-fold cover the $3$-torus $M_{1}$. From the isometries at [CM22, page 3], the covering group $C_{6}$ acting on some torus $S$ (chosen later) results in four copies of $S$ each representing a different homology class in $H_{2}(\mathbb{T}^{3};\mathbb{Z})$. Denote the projection by $\pi_{6}:M_{1}\rightarrow M_{6}$. Take $S$ to be a torus with normal vector $(1,1,1)$, then the other the other tori have normals (1,1,1),(1,1,1)(-1,-1,1),(-1,1,-1), and (1,1,1)(1,-1,-1), resulting in contractible complementary regions by 4.3. Thus $\bar{S}\coloneqq\pi_{6}(S)$ is a filling surface in $M_{6}$.
@@ 4.2 The measures and ranks of a filling surface and its ambient manifold
Suppose $[S]$ is filling. By 2.7, a least area representative $S$ cuts $M$ into balls. The boundary components of MSM-S together form two copies of $S$. By [Rat19, (3.4.6)], the hyperbolic metric in spherical coordinates is
@@
(4.1) $\textnormal{Vol}(B(r))=\textnormal{Area}(S^{2})\int_{0}^{r}\sinh^{2}t\,\,\mathrm{d}t=\frac{\pi}{2}(e^{2r}-e^{-2r}-4r)$
@@
(4.2) $\textnormal{Area}(\partial B(r))=\textnormal{Area}(S^{2})\sinh^{2}r=\pi(e^{2r}+e^{-2r}-2).$
@@
$\textnormal{Area}(M-S)=2\textnormal{Area}(S)\geq\textnormal{Area}(\partial B(r))\geq 2\textnormal{Vol}(B(r))=2\textnormal{Vol}(M).$
Equality is only achieved at $r=0$ and hence it is a strict inequality. Let $V(r)=\textnormal{Vol}(B(r))$. [BGLS10, p. 2204] gives an explicit formula relating the rank of a lattice to its covolume, which for $PSL_{2}(\mathbb{C})$ is
$d(\Gamma)\leq\frac{V(1.25\epsilon)}{V(0.25\epsilon)^{2}}\textnormal{Vol}(M)<1.698\times 10^{9}\textnormal{Vol}(M),$

### v3

@@ 3.6 Least area non-Fuchsian surfaces fail to minimize intersection numbers
Let $M$ be a hyperbolic $3$-manifold. Then for any least area non-Fuchsian surface $S$, there exists a closed geodesic $\gamma$ such that the intersection number $|S\cap\gamma|$ is strictly larger than the geometric intersection number $S]\cdot[\gamma]$.
@@
Take any non Fuchsian least area surface $S$ of genus $g$ in $M$. Since minimal surfaces in $3$-manifolds are smoothly immersed by 3.12 and non transverse self-intersections are isolated by [FHS83, Lemma 1.4], there is a point $p\in S$ not in the self-intersection locus of $S$. There exists a tangent plane at $p$ and also a plane $\Pi\subset M$ tangent to $S$ at $p$.
@@
We can draw two small open balls $B_{1},B_{2}$ supported in a small neighborhood $U$ of $p$, so that $U$ is away from the self-intersection locus. See Figure 3. The set of geodesics which pass through $B_{1}$ and $B_{2}$ in $M$ contains an open subset. By [KH95, Theorem 17.6.2, Corollary 6.4.20], the collection of closed geodesics is a dense subset of the unit tangent bundle of $M$. Hence, there is a closed geodesic $\gamma$ in $M$ which passes through $B_{1}$ and $B_{2}$. Then the geodesic $\gamma$ intersects the least area surface $S$ at two points which are non essential intersections, and can be removed after applying a homotopy to $\gamma$. ∎
3.7 Cubulations and one-orbit of hyperplanes
For backgrounds and basic definitions of cube complex and its connection to $3$-manifolds, we refer the readers to [AFW15, FH21, BW12] and the references therein.
A cube complex is a metric polyhedral complex all of whose cells are unit cubes, i.e., it is the quotient of a disjoint union of copies of unit cubes under an equivalence relation generated by a set of isometric identifications of faces of cubes. Throughout this subsection, let $X$ be a CAT(0) cube complex. We define an equivalence relation on the edges of $X$ generated by identifying opposite edges of some square in $X$. Given an equivalence class $[e]$ of edges, the hyperplane $w$ dual to $[e]$ is the collection of midcubes which intersect edges in $[e]$. The complement XwX-w consists of two halfspaces.
Let $G$ be a finitely generated group with a Cayley graph $\Delta$. A subgroup $H\leq G$ is codimension-$1$ if it has a finite neighborhood $N_{r}(H)$ such that $\Delta-N_{r}(H)$ contains at least two $H$-orbits of deep components, which do not lie in any $N_{s}(H)$. For $M$ a closed hyperbolic $3$-manifold and $G=\pi_{1}(M)$, a quasi-Fuchsian surface subgroup of $G$ is codimension-$1$ and quasiconvex.
A CAT(0) cube complex $X$ is essential, if for each hyperplane $w$, each of the associated halfspaces contains points in $X$ arbitrarily far from $w$. If a group $G$ acts by automorphism on $X$, the action is essential if for any point $x$ in the zero-skeleton of $X$, and each hyperplane $w$, each of the associated halfspaces contains points in $G\cdot x$ arbitrarily far from $w$.
The cube complex $X$ is hyperplane-essential if each hyperplane $w$, regarded itself as a CAT(0) cube complex, is essential. The action of $G$ on $X$ is hyperplane-essential if each hyperplane $w$ has the property that the stabilizer of $w$ acts essentially on $w$.
If $G$ is a finitely generated group with a finite collection of codimension-$1$ subgroups $H_{1},\cdots,H_{k}$, Sageev [Sag95] introduced a powerful construction that produces an action of $G$ on a CAT(0) cube complex $X$ that is dual to a system of walls corresponding to these subgroups. This elucidates many structures of the group. Moreover, he established the following useful criterion for compactness property.
Theorem 3.22 (Sageev, Sag97).
Let $G$ be a word-hyperbolic group, and $H_{1},...,H_{k}$ be a collection of quasiconvex codimension-$1$ subgroups. Then the action of $G$ on the dual cube complex is cocompact.
Proof of 1.8.
Proof.
Since $M$ is closed hyperbolic, $\pi_{1}(M)$ is word-hyperbolic. Since a strongly filling quasi-Fuchsian surface (or totally geodesic hypersurface) $[S]$ separates any two points on the Gromov boundary of $G$, the singleton $\{[S]\}$ satisfies the assumption of [BW12, Theorem 1.4]. Thus $G$ acts properly and cocompactly on the dual CAT(0) cube complex $X$. Since $[S]$ is quasi-convex codimension-$1$, it follows that $X$ is essential.
The orbits of the hyperplanes correspond to the image $H$ under the natural action by $\pi_{1}(M)$. The stabilizers of the hyperplanes correspond to conjugation of $H\cong\pi_{1}(S)$. Since $S$ is compact, $\pi_{1}(S)$ is again a word-hyperbolic group. With the existence of filling closed geodesics, using similar reasoning above, we deduce that $X$ is hyperplane-essential.
Moreover, $[S]$ strongly filling corresponds to the dual cube complex having one-orbit of hyperplanes.
Since non commensurable strongly filling surfaces have different limit sets on $\partial G$, they give rise to cubulations that are not $\pi_{1}(M)$-equivariantly isometric. As a consequence of Theorem 1.1 and Theorem 1.6, we establish 1.8. ∎
This partially answers a question of Fioravanti-Hagen on [FH21, p. 879]. For any Gromov-hyperbolic group $G$ that admits a proper, cocompact action on a CAT(0) cube complex, Fioravanti-Hagen [FH21] establish the existence of a CAT(0) cube complex $X$ on which $G$ acts properly, cocompactly, and essentially with a single orbit of hyperplanes, or acts properly, cocompactly, essentially, and hyperplane-essentially. Moreover, if we apply their theorems to the case where $G=\pi_{1}(M)$ for $M$ a closed hyperbolic $3$-manifold, it is not clear that their hyperplane-stabilizer is a surface subgroup [^1] M. Hagen informed the author that their construction applied to a filling collection of Fuchsian surfaces should give hyperplane-stabilizers that are surface subgroups.. This is because their cubulation comes from taking a filling collection $A$ of finitely many surfaces constructed by [KM12b] and then applying the boundary criterion [BW12]. Surfaces in $A$ may coincide on pants instead of just intersecting at closed curves. Thus it is not clear their hyperplane-stabilizers are surface subgroups. In Kahn-Wright’s construction [KW21], one uses two copies of each good pants. See also 3.18.
Remark 3.23 (An analogy between Fioravanti-Hagen’s remark and our results on ubiquity).
Fioravanti-Hagen remark on [FH21, p.880] that for certain $G$ and $X$ (including our cases), an arbitrarily small deformation of one of the original walls suffices to obtain a proper action with a single orbit of hyperplanes. In particular, asymptotically geodesic surfaces (resp. totally geodesic hypersurfaces) are ubiquitous by 1.9 and eventually strongly filling Theorems 1.1 and 1.6; thus in an arbitrarily small neighborhood of a geometric circle (resp. geometric sphere $S^{n-2}$) at $\partial\widetilde{M}$, there exists a limit set of a strongly filling surface (resp. hypersurface). The technique of [FH21] is purely combinatorial.
@@ 4 From filling surfaces to the topology and geometry of the ambient manifold
Dunfield first suggested the existence of a filling $\pi_{1}$-injective surfaces in non hyperbolic manifolds, such as the Hantzsche-Wendt manifold. In this section, we will classify closed orientable geometric $3$-manifolds which contain a filling $\pi_{1}$-injective surface (Theorem 1.10) and prove the isoperimetric inequality between a filling surface and $M$ (Theorem 1.11).
@@
Let $M$ be a Seifert fiber space with a geometric structure. Let $S$ be a closed surface not equal to $S^{2},\mathbb{R}P^{2}$, and $f:S\rightarrow M$ a minimal $\pi_{1}$-injective immersion. Then $F$ is either vertical or horizontal.
@@
Scott [Sco78] shows that the fundamental group of a Seifert fibered manifold $M$ is LERF. Thus to understand the complementary regions to an immersed $\pi_{1}$-injective surface $S$ in $M$, we can pass to a finite cover $N$ of $M$ so that the preimage of $S$ in $N$ is a finite union of embedded $\pi_{1}$-injective surfaces that are invariant under the covering group $C$. 3.5 implies that it suffices to show that no such collection of embedded $\pi_{1}$-injective surfaces fills $N$.
A closed $3$-manifold with $S^{2}\times\mathbb{R}$ geometry is finitely covered by $S^{2}\times S^{1}$. Then $S^{2}\times\theta$ is the only $\pi_{1}$-injective surface up to homotopy, and any finite union of these surfaces is non-filling.
@@
    the Hantzsche–Wendt manifold, whose fundamental domain (in the lower right-hand corner of Figure 4) is made from two unit cubes whose faces are glued according to the markings (the unmarked top and bottom are glued with translations).
@@
The $3$-torus has no filling surfaces, since every horizontal surface is a torus whose complement is $\mathbb{T}^{2}\times I$. Using ideas in 3.6 and 3.2, we prove
@@
Readers can compare 4.3 with 3.6.
For $i=2,3,6$, $C_{i}$ denotes the covering groups corresponding to the covering $M_{1}\rightarrow M_{i}$. For $i=4,5$, $C_{i}$ refer to the groups corresponding to the covering from hexagonal $\mathbb{T}^{3}$ (the hexagonal fundamental domain with opposite faces glued to each other) to $M_{i}$.
@@
The last one is the Hantzsche-Wendt manifold $M_{6}$, which has as a $4$-fold cover the $3$-torus $M_{1}$. From the isometries at [CM22, page 3], the covering group $C_{6}$ acting on some torus $S$ (chosen later) results in four copies of $S$ each representing a different homology class in $H_{2}(\mathbb{T}^{3};\mathbb{Z})$. Denote the projection by $\pi_{6}:M_{1}\rightarrow M_{6}$. Take $S$ to be a torus with normal vector $(1,1,1)$, then the other the other tori have normals (1,1,1),(1,1,1)(-1,-1,1),(-1,1,-1), and (1,1,1)(1,-1,-1), resulting in contractible complementary regions by 4.3. Thus $\bar{S}\coloneqq\pi_{6}(S)$ is a filling surface in $M_{6}$.
@@ 4.2 The measures and ranks of a filling surface and its ambient manifold
Let $\textnormal{rank}(\cdot)$ denote the rank of a group (the minimum number of generators). Suppose $[S]$ is filling. By 2.7, a least area representative $S$ cuts $M$ into balls. Moreover, $\textnormal{Area}(S)\leq 4\pi(g-1)$ and $\textnormal{rank}(\pi_{1}(S))=2g$. The boundary components of MSM-S together form two copies of $S$. By [Rat19, (3.4.6)], the hyperbolic metric in spherical coordinates is
@@
$\textnormal{Vol}(B(r))=\textnormal{Area}(S^{2})\int_{0}^{r}\sinh^{2}t\,\,\mathrm{d}t=\frac{\pi}{2}(e^{2r}-e^{-2r}-4r)$ (4.1)
@@
$\textnormal{Area}(\partial B(r))=\textnormal{Area}(S^{2})\sinh^{2}r=\pi(e^{2r}+e^{-2r}-2).$ (4.2)
@@
$\textnormal{Area}(\partial(M-S))=2\textnormal{Area}(S)\geq\textnormal{Area}(\partial B(r))\geq 2\textnormal{Vol}(B(r))=2\textnormal{Vol}(M).$
Equality is only achieved at $r=0$ and hence it is a strict inequality. Let $V(r)=\textnormal{Vol}(B(r))$ and $\textnormal{rank}(\pi_{1}(M))$ the rank of $\pi_{1}(M)$. [BGLS10, p. 2204] gives an explicit formula relating the rank of a lattice to its covolume, which for $PSL_{2}(\mathbb{C})$ is
$\textnormal{rank}(\pi_{1}(M))\leq\frac{V(1.25\epsilon)}{V(0.25\epsilon)^{2}}\textnormal{Vol}(M)<1.698\times 10^{9}\textnormal{Vol}(M),$